OpenStudy (anonymous):

PLEASE HELP!! I tried everything. I got answers $18.20 & $20 Chris Nazarro had total travel expenses of $872. His transportation cost was $136. The hotel cost was $210 and business materials cost $500. If his lunch cost 30% of his dinner, how much was dinner to the nearest cent?

OpenStudy (anonymous):

Let x be the cost of dinner. Then 0.3x is the cost of lunch. Add up all the costs and set them equal to 872 136 + 210 + 500 + x + 0.3x = 872

OpenStudy (anonymous):

846+ x + 0.3x=872

OpenStudy (anonymous):

x+0.3x=26

OpenStudy (anonymous):

so will both xs be added together making it 1.3?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

then i divide that?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

I get $20 /: thats how I did it the first time.

OpenStudy (anonymous):

that's what I got. what are you supposed to get?

OpenStudy (anonymous):

Oh wait. I just put in the answer as $20.00 and got it right i forgot the .00 thank you so much!

OpenStudy (anonymous):

you're welcome